# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v1

@@ III Numerical test
In this section, we show a proof-of-principle demonstration of our analog simulation approach using the one-dimensional, two-electron Shin-Metiu model [56, 57] and its implementation on a trapped ion quantum computer. The model consists of two electrons, two fixed ions, and one moving ion between the fixed ions displaced by a distance $L=5.4$ a.u. (Fig. 3a), where the origin is set to the middle point of the two fixed ions. The computational details for all simulations performed in this section, including the parameters for the model Hamiltonian, are given in the Supplemental Material and Refs. 56, 57, 58, 14.
@@
We construct the electronic Fock space with four spin orbitals $\{|\phi_{i}\rangle\}$ using the spin-up and spin-down configurations of two diabatic spatial orbitals, $\eta_{a}$ and $\eta_{b}$ in Fig. 3c and 2d, respectively, i.e. $\varphi_{1}=\varphi_{2}=\eta_{a}$, and $\varphi_{3}=\varphi_{4}=\eta_{b}$. As a result, a total of four ions ($max(N_{\mathrm{mode}}=1,N_{\mathrm{o}}=4)$) are needed in the experimental trapped ion setup for the simulation. The diabatic orbitals were obtained by numerically integrating two adiabatic orbitals from the one-electron Shin-Metiu Hamiltonian [56] subject to the diabatization condition. The resulting orbitals are localized around the left and right fixed ions, but delocalize slightly as the moving ion approaches the fixed ion (Fig. 3c and 3d). We replaced the Coulomb potential for $V_{\mathrm{nn}}$ with a harmonic potential to simplify the Hamiltonian and to confine the spatial extent of the wavefunction. The one- and two-electron integrals are truncated at first order in the nuclear position: $v(\hat{Q})\approx v_{0}+v_{1}\hat{Q}$, where $\hat{Q}=(\hat{b}^{\dagger}+\hat{b})/\sqrt{2\omega}$ with ladder operators $\hat{b}^{\dagger}$ and $\hat{b}$ of our single mode with frequency $\omega$ of the harmonic nuclear repulsion potential. With these approximations, the mapped Hamiltonian can be written as
$\hat{H}_{\mathrm{cMQB}}=\omega\hat{b}^{\dagger}\hat{b}+\sum_{I}(V_{0}^{I}+V_{1}^{I}(\hat{b}^{\dagger}+\hat{b}))\hat{P}_{1}^{I}\hat{P}_{2}^{I}\hat{P}_{3}^{I}\hat{P}_{4}^{I}.$ (9)
@@
The model system has only a single nuclear degree of freedom, meaning the normal mode coordinate is equal to the position of the moving ion multiplied by the square root of the mass of the ion $M$, i.e. $Q=\sqrt{M}R$. We show results in terms of the unweighted coordinate $R$ rather than $Q$ to show the ion and electrons in the same position space.
@@
We initialized the molecular state as $|\Psi(0)\rangle=\hat{D}(R_{0}\sqrt{M\omega/2})|0\rangle_{\mathrm{n}}\otimes|1100\rangle_{\mathrm{q}}$ (Fig. 3d) , where the ground state wavefunction of the harmonic oscillator for bosonic degrees of freedom is displaced by $R_{0}=0.1$ a.u. in real space, and electrons are in the closed shell configuration of the orbital $\eta_{a}$ (Fig. 3b). To achieve the initial state experimentally on an ion trap, we would first prepare the ground state of two bosonic modes and of each of the qubits, $|0000\rangle_{\mathrm{q}}$. The additional bosonic mode is required to implement qubit entanglement operations. Then, we would apply a digital quantum operation $\hat{X}_{1}\hat{X}_{2}$, and subsequently apply the displacement operator using spin-dependent force on ions with a laser interaction [59].
@@
For the time-evolution, we assume that the base Hamiltonian of the ion trap (the first term in Eq. 9) is always present during the simulation and rescale the Trotter steps to compensate for it [39]. The Trotterized cMQB time-evolution operators could be achieved by either one of methods introduced in Sec. II.1.2. To achieve the first approach, we would apply a series of digital quantum gates on the trapped ions with single-ion addressing. The second bosonic mode would be used to apply the entangling gates [59].
@@
Figure 3: Results for theoretical simulations of the one-dimensional, two-electron Shin-Metiu model with exact time-evolution, cMQB time-evolution, and the GBOA. a, The model system consists of two fixed ions (white circles) displaced by distance $L$, a moving ion (grey circle) at a position $R$, and two electrons (green circles) at positions $r_{1}$ and $r_{2}$. The origin is the average of the fixed ion positions. The diabatic orbitals: b, $\eta_{a}(r;R)$ localized around the left fixed ion, and c, $\eta_{b}(r;R)$ localized around the right fixed ion. The density functions $\rho(r,R,t)$ at $t=56.1$ a.u. (eg) and $t=1514.4$ a.u. (eg), with exact time-evolution (e and h), cMQB time-evolution ($\Delta t=5.6$ a.u.) (f and h), and GBOA (g and j). Spatial functions in bj are normalized to their maximum values.
@@
In Fig. 3, we show the density functions $\rho(\mathbf{r},\mathbf{Q},t)$ obtained from exact time-evolution, Trotterized cMQB time-evolution, and the GBOA. Experimentally, determination of the density function would require an ancilla qubit which controls the operator for the density measurement. Controlled Pauli strings correspond to CNOT and single-qubit gates, and controlled displacement operators can be achieved with a spin-dependent force on the trapped ions [59, 55]. The sum of all orbital pairs with Fourier transforms over a grid of displacements yields the density, as shown in Eq. 6. For this example, only 12 sets of measurements are required: two Pauli strings with four on-diagonal and two non-zero off-diagonal pairs. This number of measurements can be halved because the densities of the two spin components are equal. We can obtain a nuclear density resolution of 0.02 a.u. using a grid spacing of 1.26 a.u. in the momentum space with 250 points.
The Trotterized cMQB time-evolution reproduces the exact density with converged Trotter step $\Delta t=5.6$ a.u (Fig. 3f,i), where the fidelity of the corresponding wavefunction is reported in the Supplemental Material. At $t=1514.4$ a.u., the non-zero densities on both sides of the fixed ions (Fig. 3h,i) indicates that the moving ion induces electron transfer from the left to the right fixed ions via vibronic coupling.
To highlight inaccuracies that originate from the GBOA, we perform equivalent dynamics simulations in the BO framework (Eq. 8) with a truncated BO basis. In this simulation, we prepare the same initial state as the pre-BO simulation (Fig. 3d). For two electrons in two spatial orbitals with a singlet spin multiplicity, $N_{\mathrm{CSF}}=3$. The ground BO state is excluded from the dynamics for the GBOA ($N_{\mathrm{BO}}=2$), which has relatively small NACs with the first excited state.
@@
The density functions obtained with the GBOA are shown in Figs. 3g,j. Within a short simulation time (56.1 a.u.), there is a spurious electron transfer from the left fixed ion to the right (Fig. 3g). This electron transfer does not occur in the exact limit (Fig. 3e) and the Trotterized cMQB time-evolution (Fig. 3f) due to vibronic coupling to the ground BO state. As a result, the subsequent density deviates even further, as seen in Fig. 3j. This demonstrates how the GBOA may result in inaccurate molecular vibronic dynamics when a molecule reaches an unanticipated NAC region coupled to a BO state outside of the truncated BO basis. Our method spans the FCI space without basis set truncation with a given orbital basis set, so there is no need to pre-determine the truncation of electronic states.
IV Discussion
@@
Figure 4: Resource and computational cost comparison of our cMQB approach with classical and quantum algorithms demonstrating the advantages of cMQB over existing methods. a, Resource scaling of electronic degrees of freedom with respect to $N_{\mathrm{o}}$ in a logarithmic scale for classical ($\log_{2}(N_{\mathrm{CSF}})$) and quantum pre-BO simulations ($\log_{2}(N_{\mathrm{q}})$) with singlet electronic states. The resource scaling of our method represented by the red line shows smaller scaling than those of equivalent classical simulations represented by the blue lines for the maximum (dashed), minimum (dotted), and $N_{\mathrm{o}}/N_{\mathrm{e}}=6$ case (solid) for $N_{\mathrm{o}}>6$. b, Scaling in the number of interactions/gates required to implement the time-evolution operator ($\log(N_{\mathrm{op}})$) with respect to $N_{\mathrm{o}}$ for our cMQB approach (red), and existing quantum algorithms in the BO framework (yellow) labeled as "qBO" for the maximum (dashed), minimum (dotted), and $N_{\mathrm{o}}/N_{\mathrm{e}}=6$ case (solid). The cMQB approach requires smaller scaling than qBO for $N_{\mathrm{o}}>6$. c, The number of qubits ($N_{\mathrm{q}}$) required vs. the number of electrons $N_{\mathrm{e}}$ for the pre-BO quantum algorithm proposed in Ref. 34 using grids where $N_{\mathrm{q}}=3nN_{\mathrm{e}}$ with $n=10$ (black) and our cMQB approach with $N_{\mathrm{q}}=6N_{\mathrm{e}}$ (red). Our cMQB method requires smaller number of qubits compared to the previous pre-BO quantum algorithm.
@@
We first compare the resource and computational costs of our cMQB pre-BO analog simulation approach with those of existing classical and quantum algorithms, where the advantages of our approach are highlighted in Fig. 4.
@@
Here, we discuss singlet states ($S=0$) since most ground-state electron configurations of neutral organic molecules are singlets since they have an even number of electrons and a lack of orbital degeneracy. The number of singlet CSFs is minimum ($O(N_{\mathrm{o}}^{2})$) when $N_{\mathrm{e}}=2$ or $N_{\mathrm{e}}=N_{\mathrm{o}}-2$, and maximum ($O(2^{N_{\mathrm{o}}}/N_{\mathrm{o}}^{2})$) when $2N_{\mathrm{e}}=N_{\mathrm{o}}$ for a fixed $N_{\mathrm{o}}$. Practical simulations fall in between these two ranges with a number of orbitals roughly proportional to the number of electrons, with the ratio depending on the basis set. For example, conjugated polyene chains with a double-zeta basis set have $N_{\mathrm{o}}/N_{\mathrm{e}}=6$. In general, the classical resources scale as $O(N_{\mathrm{o}}^{-2}(m(m-1)^{1/m-1})^{N_{\mathrm{o}}})$ for $N_{\mathrm{o}}=mN_{\mathrm{e}}$ ($m>1$) at large $N_{\mathrm{o}}$ limit. In Fig. 4a, we show classical resource scaling for singlet states in comparison with our approach. The linear scaling of our approach has a clear advantage over classical methods, even over the minimum $N_{\mathrm{CSF}}$ case for $N_{\mathrm{o}}>6$.
@@
For the nuclear degrees of freedom, the classical simulation has an exponential resource scaling of $O(\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}})$ where $N_{\nu,\mathrm{bas}}$ is the size of basis for mode $\nu$. Our enconding on cMQB simulators has a linear scaling $O(N_{\mathrm{mode}})$, which is equal to that of an MQB mapping [39].
Like classical methods, the number of qubits can be reduced by using a complete active space. The active space defines a subset of orbitals for which the exact solution is found, while the remaining (inactive) orbitals have fixed occupations [60]. Only the active orbitals need to be encoded in our simulation because the terms in the Hamiltonian corresponding to inactive orbitals are functions of nuclear coordinates, meaning their effect can be included in the electron integral terms of active orbitals. Likewise, strongly confined modes can be approximately excluded from the nuclear part of the wavefunction for a reduced-dimensional simulation. Together, these reductions in the basis size allow for chemically relevant simulations on near-term quantum hardware.
@@
The simulation cost, which can be measured by the number of interaction terms required, scales as $O(N_{\mathrm{o}}^{4}N^{k}_{\mathrm{mode}})$ where $k$ is the maximum order of the Taylor expansion for the functions of bosonic degrees of freedom $f_{I}$. This cost also remains far smaller than the exponential cost of quantum dynamics simulations on classical computers: $O(N_{\mathrm{CSF}}^{2}\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}}^{2})$.
@@
Previous quantum algorithms, including the MQB approach, can potentially have a smaller quantum resource scaling for electronic degrees of freedom than our cMQB method: $\log_{2}(N_{\mathrm{CSF}})<N_{\mathrm{o}}$ for the complete CSF basis [37, 38, 39] since the CSF space is a subspace of the spin orbital Fock space. However, the simulation cost $N_{\mathrm{op}}$ scales as $O(N_{\mathrm{CSF}}^{2})$ in the existing quantum simulation approaches, whereas our mapping scales as $O(N_{\mathrm{o}}^{4})$ (Fig. 4b). Furthermore, an equivalent approach would require fully diagonalizing the FCI Hamiltonian on a classical computer at a prohibitive cost ($O(N_{\mathrm{CSF}}^{3})$), whereas our approach requires only the pre-calculation of electron integrals that scales as $O(N_{\mathrm{o}}^{4})$. The GBOA is valid approximation in many photo-initiated reactions provided that the photon energy absorbed by a system is known so that we can correctly reduce the number of BO states according to that energy window. However, the GBOA breaks down in strong field processes due to the high energy and strong correlation. Furthermore, even if the number of BO states can be reduced, the BO states are obtained by approximate For the nuclear degrees of freedom, our approach has an advantage over other digital or hybrid quantum algorithms for molecular quantum dynamics [37, 38] since our approach has the same resource scaling as the MQB mapping [39] due to the direct mapping on the bosonic modes.
@@
The first proposal of a quantum algorithm for the pre-BO molecular dynamics used a real-space, first quantization representation which has a linear resource scaling with respect to the number of grid points per each degrees of freedom: $nN_{\mathrm{mode}}+3nN_{\mathrm{e}}$, where $n$ is the number of qubits required for each degree of freedom. The authors suggested a minimum grid size of $n=10$ [34]. Although their approach also scales linearly with the number of electrons, our approach has a lower resource cost for most molecules with a reasonable basis set size, where $N_{\mathrm{mode}}<N_{\mathrm{e}}$ and $N_{\mathrm{o}}<30N_{\mathrm{e}}$ (Fig. 4c).
@@
A key approximation in our approach is the exclusion of the Coriolis (rotational-vibrational) coupling from the Hamiltonian. This approximation is not only useful for simplifying the Hamiltonian, but it also comes with an advantage for physical observables. Any observable quantity that depends on the molecular (Eckart) frame is averaged out over the rotational wavefunction, including electronic and nuclear densities. Experimental measurement of these observables in a lab frame results from spontaneous symmetry breaking between the molecule of interest and its environment [61]. This gives a potential advantage for interpretability of our wavefunction compared to other quantum simulation proposals [34, 36]. Furthermore, on the timescale of ultrafast processes (fs–ps), we expect that the effect of Coriolis coupling would be a negligible perturbation on the vibronic states for due to the relatively small rotational frequencies. Future work could explore how this perturbation could be included as an environmental effect without compromising the wavefunction interpretability.
@@
An additional approximation is the lack of relativistic effects, which becomes particularly relevant for heavier atomic nuclei. Classical simulations need to consider additional CSFs with different spin multiplicities to include relativistic effects such as spin-orbit coupling. Therefore the classical resource scaling becomes even larger if intersystem crossing between different multiplicities is included. In contrast, since our pre-BO ansatz spans the entire electronic Fock space for a given spin orbital basis, our approach can describe intersystem crossing by adding spin-orbit coupling terms to the Hamiltonian, without any additional resources required. Based on the form of the vibronic Hamiltonian, our Hamiltonian can be considered as an extended version of electron-phonon coupling models used in the solid state physics field such as the Hubbard-Holstein model [62]. Therefore, our approach can be readily extended to dynamics involving intersystem crossing and translated onto electron-phonon coupling dynamics in solids.
As an analog simulation approach, our approach is subject to environmental noise over the course of the simulation. The noise does not equally affect electronic and vibrational degrees of freedom, since they are encoded in different resources of the simulator. However, because cMQB simulators use bosonic modes to achieve entangling gates, the electronic degrees of freedom are subject to noise on qubit states and bosonic modes. Nonetheless, previous work [39, 63, 64] has shown that controlled noise can be used as an advantage for analog simulators by simulating open quantum systems with a minimal cost. In future work, we will explore noise effects to determine the threshold for a near-term quantum advantage in quantum chemistry.
V Conclusion

### v2

@@ II.4 Numerical test
We show a proof-of-principle demonstration of our analog simulation approach using the one-dimensional, two-electron Shin-Metiu model [62, 63] and its implementation on a trapped ion quantum simulator, which has been used for the quantum simulation of chemical dynamics for several systems [41, 43, 44, 64]. The model consists of two electrons, two fixed ions, and one moving ion between the fixed ions displaced by a distance $L=5.4$ a.u. [Fig. 3(a)], where the origin is set to the middle point of the two fixed ions. The computational details for all simulations performed in this section, including the parameters for the model Hamiltonian, are given in the Supplemental Material [46].
@@
We construct the electronic Fock space with four spin orbitals $\{|\phi_{i}\rangle\}$ using the spin-up and spin-down configurations of two diabatic spatial orbitals, $\eta_{a}$ and $\eta_{b}$ in Figs. 3(b) and 3(c) respectively, i.e. $\varphi_{1}=\varphi_{3}=\eta_{a}$, and $\varphi_{2}=\varphi_{4}=\eta_{b}$. As a result, a total of four ions are needed in the experimental trapped ion setup for the simulation, with only two of the 12 available motional modes needed for the simulation. The diabatic orbitals were obtained by numerically integrating two adiabatic orbitals from the one-electron Shin-Metiu Hamiltonian [62] subject to the diabatization condition. The resulting orbitals are localized around the left and right fixed ions, but delocalize slightly as the moving ion approaches the fixed ion [Figs. 3(b) and 3(c)]. We replaced the Coulomb potential for $V_{\mathrm{nn}}$ with a harmonic potential to simplify the Hamiltonian and to confine the spatial extent of the wavefunction. The one- and two-electron integrals are truncated at first order in the nuclear position: $v(\hat{Q})\approx v_{0}+v_{1}\hat{Q}$, where $\hat{Q}=(\hat{b}^{\dagger}+\hat{b})/\sqrt{2\omega}$ with ladder operators $\hat{b}^{\dagger}$ and $\hat{b}$ of our single mode with frequency $\omega$ of the harmonic nuclear repulsion potential. With these approximations, the cMQB Hamiltonian can be written as
$\hat{H}_{\mathrm{cMQB}}=\omega\hat{b}^{\dagger}\hat{b}+\sum_{I}(V_{0}^{I}+V_{1}^{I}(\hat{b}^{\dagger}+\hat{b}))\hat{P}_{1}^{I}\hat{P}_{2}^{I}\hat{P}_{3}^{I}\hat{P}_{4}^{I},$ (12)
@@
where $V_{k}^{I}$ represents the effective $k$-th order coefficient of the nuclear position coupled to the $I$-th Pauli string. The model system has only a single nuclear degree of freedom, meaning the normal mode coordinate is equal to the position of the moving ion multiplied by the square root of the mass of the ion $M$, i.e. $Q=\sqrt{M}R$. We show results in terms of the unweighted coordinate $R$ rather than $Q$ to show the ion and electrons in the same position space.
@@
We consider the initial molecular state $|\Psi(0)\rangle=\hat{D}(R_{0}\sqrt{M\omega/2})|0\rangle_{\mathrm{n}}\otimes|1010\rangle_{\mathrm{q}}$ [Fig. 3(d)] , where the ground state wavefunction of the harmonic oscillator for bosonic degrees of freedom is displaced by $R_{0}=0.1$ a.u. in real space, and electrons are in the closed shell configuration of the orbital $\eta_{a}$ [Fig. 3(b)]. To achieve the initial state experimentally on an ion trap, we would first prepare the ground state of two bosonic modes and of each of the qubits, $|0000\rangle_{\mathrm{q}}$. Then, we would apply a digital quantum operation $\hat{X}_{1}\hat{X}_{3}$, and subsequently apply the displacement operator using a spin-dependent force on the ions with a laser [65]. During the simulation, an additional bosonic mode is required to implement qubit entanglement operations.
@@
For the time-evolution, we assume that the base Hamiltonian of the ion trap (the first term in Eq. (12)) is always present during the simulation and rescale the Trotter steps to compensate for it [41]. To achieve the digital-analog time evolution, we would apply a series of digital quantum gates on the trapped ions with single-ion addressing. The second bosonic mode would be used to apply the entangling gates [65].
@@
Figure 3: Results for theoretical simulations of the one-dimensional, two-electron Shin-Metiu model with exact time-evolution, our cMQB approach, and the GBOA. (a) The model system consists of two fixed ions (white circles) displaced by distance $L$, a moving ion (grey circle) at a position $R$, and two electrons (green circles) at positions $r_{1}$ and $r_{2}$. The origin is the average of the fixed ion positions. The diabatic orbitals are (b) $\eta_{a}(r;R)$ localized around the left fixed ion, and (c) $\eta_{b}(r;R)$ localized around the right fixed ion. The density functions $\rho(r,R,t)$ at, (d) $t=0$ a.u., (e)–(g) $t=56.1$ a.u., and (h)–(j) $t=1514.4$ a.u., with exact time-evolution ((e) and (h)), Trotterized cMQB time-evolution ($\Delta t=5.6$ a.u.) ((f) and (i)), and GBOA ((g) and (j)). Spatial functions in (b)–(j) are normalized to their maximum values.
@@
In Fig. 3, we show the density functions $\rho(r,R,t)$ obtained from exact time-evolution, Trotterized cMQB time-evolution, and the GBOA. Experimental determination of the density function would require an ancilla qubit which controls the operator for the density measurement. Controlled Pauli strings correspond to CNOT and single-qubit gates, and controlled displacement operators can be achieved with a spin-dependent force on the trapped ions [65, 59]. The sum of all orbital pairs with Fourier transforms over a grid of displacements yields the density, as previously shown in Eq. (6). For this example, only 12 sets of measurements are required: two Pauli strings with four on-diagonal and two non-zero off-diagonal pairs. This number of measurements can be halved because the densities of the two spin components are equal. We can obtain a nuclear density resolution of 0.02 a.u. using a grid spacing of 1.26 a.u. in the momentum space with 250 points.
@@
$N_{\mathrm{CSF}}=3$ is the size of complete electronic basis of singlet electronic states with two electrons in two spatial orbitals. We have excluded the ground BO state from the dynamics with the GBOA ($N_{\mathrm{BO}}=2$), which has relatively small NACs with the first excited state (see Supplemental Material [46] for densities of electronic states in the BO framework). Within a short simulation time (56.1 a.u.), the GBOA density shows a spurious electron transfer from the left fixed ion to the right [Fig. 3(g)] where the subsequent density deviates even further [Fig. 3(j)]. The electron transfer at $t=56.1$ a.u. does not occur in the exact simulation [Fig. 3(e)] due to vibronic coupling to the ground BO state. This demonstrates how the GBOA may result in inaccurate molecular vibronic dynamics when a molecule reaches an unanticipated NAC region coupled to a BO state outside of the truncated BO basis. On the other hand, the Trotterized cMQB time-evolution reproduces the exact density with a converged Trotter step of $\Delta t=5.6$ a.u [Figs. 3(f) and 3(i)] with the fidelity $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}>0.95$, where the convergence of fidelity with respect to the Trotter step is reported in the Supplemental Material [46].
Figure 4: Comparison of noise effects due to motional decoherence with different Trotter step sizes for cMQB simulations. (a) Time-evolution of the fractional occupation number of $\phi_{1}$. (b) Time-evolution of the fidelity of the molecular wavefunctions: $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}$. The blue, red, and green lines represent the noise simulation results with $\Delta t=5.6$, $9.6$, and $16.8$ a.u., respectively, where the dashed lines represent the corresponding closed system simulation results and the black line represents the exact closed system simulation result without Trotterization.
We perform open quantum system dynamics simulations using Eq. (8) with Trotterization to estimate the noise effect on the dynamics of our model system, using a Rabi rate for ion-laser interaction of $\Omega=2\pi\times 1.0$ MHz [66] and the native vibrational decoherence rate of $\gamma_{\mathrm{mot}}^{\mathrm{d}}=30$ s-1 [61] based on existing ion trap cMQB devices. We report the time evolution of the fractional occupation numbers (FONs) of $\phi_{1}$ and the fidelities in Fig. 4 with vibrational decoherence and the indirect qubit noise effect on the molecular dynamics from the motional decoherence of an ion-trap, in comparison with the result without noise. The indirect qubit noise effect is negligible compared to the direct vibrational decoherence effect on nuclear degrees of freedom in FON, while the fidelity is more sensitive (See Supplemental Material [46]).
Although the noise effect leads to a deviation from the exact dynamics, the FONs are qualitatively well reproduced [Fig. 4(a)]. Furthermore, the actual noise effect can be minimized by circuit optimization and experimental techniques to reduce the experiment time required for cMQB simulations. For example, the cMQB Hamiltonian of our model system consists of many two-qubit Pauli strings without bosonic terms, where their time-evolution operators can each be realized by a single native two-qubit entanglement operator (e.g. using $\hat{R}_{XX}(\theta)=\exp(-i\frac{\theta}{2}\hat{X}_{1}\hat{X}_{2})$) instead of nesting a single qubit exponential operator with CNOT gates.
Figure 4(b) confirms the trade-off between Trotter error the and noise effect. The fidelity with $\Delta t=9.6$ a.u. is greater than the fidelity with $\Delta t=16.8$ a.u. due to the smaller Trotter error. However, the fidelity decreases again when the Trotter step is further reduced to $\Delta t=5.6$ a.u. due to the contribution from the digital part in the molecular scaling factor [Eqs. (9) and (10)] for the noise. Therefore, the Trotter step should be chosen carefully considering both the Trotter error and the noise.
III Discussion
@@
Figure 5: Resource and computational cost comparison of our cMQB approach with other classical and quantum algorithms. (a) Resource scaling of electronic degrees of freedom with respect to $N_{\mathrm{o}}$ in a logarithmic scale for classical ($\log_{2}(N_{\mathrm{CSF}})$) and quantum pre-BO simulations ($\log_{2}(N_{\mathrm{q}})$) with singlet electronic states. The resource scaling of our method is represented by the red line, wheras those of equivalent classical simulations is represented by the blue lines for the maximum (dashed), minimum (dotted), and $N_{\mathrm{o}}/N_{\mathrm{e}}=6$ case (solid). (b) Scaling in the number of interactions/gates required to implement the time-evolution operator ($\log(N_{\mathrm{op}})$) with respect to $N_{\mathrm{o}}$ for our approach (red), and existing quantum algorithms in the BO framework (yellow) labeled as "qBO" for the maximum (dashed), minimum (dotted), and $N_{\mathrm{o}}/N_{\mathrm{e}}=6$ case (solid). (c) The number of qubits ($N_{\mathrm{q}}$) required vs. the number of electrons $N_{\mathrm{e}}$ for the pre-BO quantum algorithm proposed in Ref. 35 using grids where $N_{\mathrm{q}}=3nN_{\mathrm{e}}$ with $n=10$ (black) and our approach with $N_{\mathrm{q}}=6N_{\mathrm{e}}$ (red).
@@
We first compare the resource and computational costs of our pre-BO analog cMQB simulation approach with those of existing classical and quantum algorithms, where the advantages of our approach are highlighted in Fig. 5.
@@
Here, we discuss singlet states ($S=0$). Most ground-state electron configurations of neutral organic molecules are singlets since they have an even number of electrons and a lack of orbital degeneracy. In general, the classical resources scale as $O(N_{\mathrm{o}}^{-2}(m(m-1)^{1/m-1})^{N_{\mathrm{o}}})$ for $N_{\mathrm{o}}=mN_{\mathrm{e}}$ ($m>1$) at the large $N_{\mathrm{o}}$ limit. The number of singlet CSFs is minimum ($O(N_{\mathrm{o}}^{2})$) when $N_{\mathrm{e}}=2$ or $N_{\mathrm{e}}=N_{\mathrm{o}}-2$, and maximum ($O(2^{N_{\mathrm{o}}}/N_{\mathrm{o}}^{2})$) when $2N_{\mathrm{e}}=N_{\mathrm{o}}$ for a fixed $N_{\mathrm{o}}$. Practical simulations fall in between these two ranges with a number of orbitals roughly proportional to the number of electrons, with the ratio depending on the basis set. For example, conjugated polyene chains with a double-zeta basis set have $N_{\mathrm{o}}/N_{\mathrm{e}}=6$. In Fig. 5(a) we show classical resource scaling for singlet states in comparison with our approach. The linear scaling of our approach has a clear advantage over classical methods, even over the minimum $N_{\mathrm{CSF}}$ case for $N_{\mathrm{o}}>6$.
@@
For the nuclear degrees of freedom, the classical simulation has an exponential resource scaling of $O(\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}})$ where $N_{\nu,\mathrm{bas}}$ is the size of basis for mode $\nu$. Our encoding on cMQB simulators has a linear scaling $O(N_{\mathrm{mode}})$, which is equal to that of an analog mixed qudit-boson (MQB) mapping [41].
@@
The simulation cost of our approach, which can be measured by the number of interaction terms required, scales as $O(N_{\mathrm{o}}^{4}N^{k}_{\mathrm{mode}})$ where $k$ is the maximum order of the Taylor expansion for the functions of bosonic degrees of freedom $f_{I}$. This cost also remains far smaller than the exponential cost of quantum dynamics simulations on classical computers: $O(N_{\mathrm{CSF}}^{2}\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}}^{2})$.
@@
Previous quantum algorithms, including the MQB approach, can potentially have a smaller quantum resource scaling for electronic degrees of freedom than our method: $\log_{2}(N_{\mathrm{CSF}})<N_{\mathrm{o}}$ for the complete CSF basis [38, 39, 41] since the CSF space is a subspace of the spin orbital Fock space. However, the simulation cost $N_{\mathrm{op}}$ would scale as $O(N_{\mathrm{CSF}}^{2})$ for electronic degrees of freedom in existing quantum simulation approaches, whereas the cMQB mapping scales as $O(N_{\mathrm{o}}^{4})$ [Fig. 5(b)]. The GBOA is a valid approximation for some photo-initiated reactions provided that the photon energy absorbed by a system is known and the BO states include all energetically accessible electronic states. Nevertheless, even if the number of BO states can be reduced, the BO states are obtained approximately since an exact pre-calculation would require fully diagonalizing the FCI Hamiltonian on a classical computer at a prohibitive cost ($O(N_{\mathrm{CSF}}^{3})$), whereas the cMQB approach requires only the pre-calculation of electron integrals that scales as $O(N_{\mathrm{o}}^{4})$. For the nuclear degrees of freedom, our approach has an advantage over other digital or hybrid quantum algorithms for molecular quantum dynamics [38, 39] since our approach has the same resource scaling as the MQB mapping [41] due to the direct mapping on the bosonic modes.
@@
The first proposal of a quantum algorithm for pre-BO molecular dynamics used a real-space, first quantization representation which has a linear resource scaling with respect to the number of grid points per each degree of freedom: $3n(N_{\mathrm{n}}+N_{\mathrm{e}})$, where $n$ is the number of qubits required for each degree of freedom and $N_{\mathrm{n}}$ represents the number of nuclei. The authors suggested a minimum grid size of $n=10$ [35]. Although their approach also scales linearly with the number of electrons, our approach has a lower resource cost for most molecules with a reasonable basis set size, where $N_{\mathrm{mode}}<N_{\mathrm{e}}$ and $N_{\mathrm{o}}<30N_{\mathrm{e}}$ [Fig. 5(c)].
The noise effect on the cMQB device depends primarily on the Trotter step size and the number of digital gates used to generate multi-qubit entanglement [Eqs. (9) and (10)]. Relative to the direct effect of motional decoherence on modes representing vibrations, the indirect qubit noise is negligible. The effect of noise is highly specific to the target problem and device; however, the scaling factors for the noise effect suggest desirable improvements in both computational and experimental aspects of the cMQB approach for a near-term quantum advantage: reducing $N_{\mathrm{CNOT}}$ via circuit optimization and reducing $t_{\mathrm{CNOT}}$ in cMQB devices.
From a different perspective, noise can be leveraged as an advantage in analog simulators by enabling the simulation of open quantum systems at minimal cost as demonstrated in previous work [67, 41, 61, 68], rather than treating it as an effect that must be suppressed.
@@
A key approximation in our approach is the exclusion of the Coriolis (rotational-vibrational) coupling from the Hamiltonian. On the timescale of ultrafast processes (fs–ps), we expect that the effect of Coriolis coupling would be a negligible perturbation on the vibronic states since the rotational frequencies are relatively small. This approximation also comes with an advantage for simulating physical observables. Any observable quantity that depends on the molecular (Eckart) frame is averaged out over the rotational wavefunction, including electronic and nuclear densities. Experimental measurement of these observables in a lab frame results from spontaneous symmetry breaking between the molecule and its environment [69]. This gives a potential advantage for interpretability of our wavefunction compared to other quantum simulation proposals [35, 37]. Future work could explore how rotations could be included as an environmental effect without compromising the wavefunction interpretability.
@@
As previously mentioned, our approach can describe intersystem crossing by considering spin-orbit coupling electron integrals and ionization by including additional terms in the Hamiltonian [Eq. (3)] without any additional resources. Such simiulations have the potential to make our approach even more advantageous since the classical and GBOA-based quantum simulations require a greater number of basis states for different spin multiplicities. Based on the form of the vibronic Hamiltonian, our Hamiltonian can be considered as an extended version of electron-phonon coupling models used in the solid state physics field such as the Hubbard-Holstein model [70]. Therefore, our approach can be readily extended to dynamics involving intersystem crossing and ionization, and translated onto electron-phonon coupling dynamics in solids.
IV Conclusion